To prepare titrations solutions, she dissolved 3.00 g of sodium thiosulfate pentahydrate in 100.0 mL of distilled water. 25.0 mL of the sodium thiosulfate solution was transferred into a 250.0 mL volumetric flask and was made up to the mark with distilled water to make solution A.

How to calculate the concentration of sodium thiosulfate?
1
Expert's answer
2020-08-28T08:07:37-0400

Weight of sodium thiosulphate = 3 g

Molecular weight of sodium thiosulphate=158.11 g/ mol

Number of mole sodium thiosulphate= 3/158.11= .019

Volume of solution= 100 ml

Concentration of solution = .019 × 1000 /100

= .19 mol/l

Take 25 ml of this solution and make 250 ml with distilled water

Now concentration is

M1V1=M2V2

.19×25=250×M2

M2= .019 mol / liter